A 100-N weight is suspended at the middle of a rope forming a letter “Y” with the angle between the ropes being 150°. Find the tensions in the rope.


1
Expert's answer
2021-05-28T07:18:20-0400

"W=2T\cdot\cos75°\to T=\frac{W}{2\cos75°}=\frac{100}{2\cdot\cos75°}=193.2\ (N)" . Answer